# ERPoT: Effective and Reliable Pose Tracking for Mobile Robots Based on Lightweight and Compact Polygon Maps

### v1

@@ V Conclusion
In this study, an efficient and reliable pose tracking framework for mobile robots called ERPoT has been proposed, which makes use of a novel prior map constructed from multiple compact polygons. Notably, this system operates solely on LiDAR data, acquiring the sparse 2D scan points through ground removal and obstacle selection. Following this, an offline map construction procedure is executed, yielding an environment map that is both lightweight and compact in its polygon representation. Ultimately, by integrating real-time 2D scan points with the prior polygon map, our approach achieves reliable long-term pose tracking based on the newly proposed point-polygon matching. Furthermore, the quantitative evaluation shows the superior performance on prior map size, pose estimation error, and runtime for pose tracking compared with the other five approaches.

### v3

@@ VI Conclusion
In this study, an efficient and reliable pose tracking framework for mobile robots called ERPoT has been proposed, which makes use of a novel prior map constructed from multiple compact polygons. Notably, this system operates solely on LiDAR data, acquiring the sparse 2D scan points through ground removal and obstacle selection. Following this, an offline map construction procedure is executed, yielding an environment map that is both lightweight and compact in its polygon representation. Ultimately, by integrating real-time 2D scan points with the prior polygon map, our approach achieves reliable long-term pose tracking based on the newly proposed point-polygon matching. Furthermore, the quantitative evaluation shows the superior performance on prior map size, pose estimation error, and runtime for pose tracking compared with the other six approaches.